OpenStudy (anonymous):

Evaluate the expression. |-7| + |3| a.)10 b.)-10 c.)4 d.)-4

OpenStudy (anonymous):

wouLd it be b ? cuss you take the hiqher numbers siqn ?

OpenStudy (anonymous):

\[|a| = a \text{, if a }\ge 0\] \[|a| = -a \text{, if a }\lt 0\]

OpenStudy (anonymous):

So in this case you have |-7| and -7 \(\lt0\) So |-7| = 7 You also have |3| and 3 \(\ge 0\) So |3| = 3

OpenStudy (anonymous):

uqhhh idont qet it ! : /

OpenStudy (anonymous):

dont you juss add the two #'s ?

OpenStudy (anonymous):

& take the larger number's siqn ?

OpenStudy (anonymous):

No.

OpenStudy (anonymous):

The absolute value makes whatever is inside it positive.

OpenStudy (anonymous):

okay so iqot 4 is that correct ?

OpenStudy (anonymous):

its 10

OpenStudy (anonymous):

why is it 1O ?

OpenStudy (anonymous):

hisbarbie, you're doing the DISTANCE. if they wnated distance, yes the answer is 4. but they want you to add the numbers.

OpenStudy (anonymous):

|-7| = 7.

OpenStudy (anonymous):

|-7| = 7 |3| = 3 |-7| + |3| = 7+3 = 10

OpenStudy (anonymous):

so |-7| + 3 = 7+3 =10

OpenStudy (anonymous):

read instructions carefully.

OpenStudy (anonymous):

ohhh . . . you quys are a qreat heLp ! : )

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

loL . can you quys check the other one iposted before this one ?

OpenStudy (anonymous):

which is it

OpenStudy (anonymous):

i c it